# ltxutilRepair the float

### v1

@@ 4.1 Assignment Strategies for Surface Codes
In this section, we describe five strategies for assigning qubits that take advantage of multiplexing and evaluate their impact on performance. These strategies assume that each photon contains a fixed number of qubits $m$ and can be applied in both scenarios (B) and (C). We assume surface code communication scenario (C), where we send the original code with $\lfloor 2d^{2}/m\rfloor$ photons.
Strategy i and ii: pair with minimum and maximum distance Strategies i and ii are applicable to the case with $m=2$. Strategy i assigns the nearest neighbor pair of qubits, which form an L-shape in the 2D lattice of the toric code, to the same photon as an example shown in Fig. 5(a). This minimizes the Manhattan distance in the lattice between the qubits in the same photon.
Strategy ii assigns the qubit at coordinates $(i,j)$ on the lattice and the qubit at coordinates $(i+d/2-1\,\textrm{mod}\,d,j+d/2-1\,\textrm{mod}\,d)$ to the same photon. An example is shown in Fig. 5. This is the arrangement that maximizes the Manhattan distance, in contrast to strategy i.
Figure 5: Examples showing possible assignments of qubits to photons. Each numbered circle denotes a qubit, and the color indicates the photon to which the qubit is assigned. Strategy i, shown in (a), minimizes the distance between qubits in the same photon, while strategy ii in (b) maximizes this distance. Note that this code is defined on the torus represented as a lattice with periodic boundary conditions.
Strategy iii: random Strategy iii is a method in which qubits are uniform-randomly selected and assigned to photons.
Strategy iv: random + threshold Strategy iv was designed to increase the separation between qubits assigned to the same photon while exploiting randomness. The strategy works by randomly selecting qubits and accepting them as the set for a photon only if the distance is greater than a certain threshold. If no suitable set of qubits can be found the threshold value is reduced.
@@
We describe the details and the motivation of each strategy in Appendix. B
@@
Figure 7: Comparison of multiplexing photon-assignment strategies for toric codes. Logical $Z$ error rate versus photon loss probability. The black curve shows the case without multiplexing. (A) The code parameters are $[\![200,2,10]\!]$ and $m=2$. The gray/brown curve shows the case for the assignment strategy for minimizing (strategy i) / maximizing (strategy ii) the distance between a pair of qubits in the same photon. The orange curve shows the case for uniformly random (strategy iii), and the blue line shows strategy iv, based on the algorithm 1. (B) The code parameters are $[\![288,2,12]\!]$ and $m=4$. $Z$ stabilizer-based assignment with light blue curve outperformed $X$ stabilizer-based assignment with light orange curve for logical $Z$ error. The mixed stabilizer-based assignment strategy performs between $X$ and $Z$. Strategy iv (blue) outperforms other assignment strategies for low error rate areas.
@@ 4.2 Performance of the assignment strategies
Fig. 7 (A) shows the performance of strategies i to iv, which are applicable to the case of $m=2$. The performance of the distance-maximizing strategy (grey) outperforms the distance-minimizing strategy (brown). Logical errors in the toric codes correspond to errors covering a longitude or meridian curve on the torus (a vertical or horizontal closed loop in the periodic lattice). When decoding erasure errors, logical errors can only occur when the qubit-support of one of these vertical or horizontal loops is entirely erased. When adjacent qubits in the lattice are erased, as in the case with the distance-minimizing photon assignment strategy, clusters of errors are more likely to cover such loops in the torus. Hence, it is not surprising that the distance-maximizing strategy outperforms the distance-minimizing strategy in our numerical simulations. It also showed that the strategies with randomness (iii and iv) outperform deterministic ones (i and ii). In particular, strategy iv outperformed the other strategies, although there was an increase in logical $Z$-error probability compared to no multiplexing.
Next, we compare the logical $Z$ error rates of strategies iii, iv, and v with $m=4$ in Fig. 7 (B). Assignment strategies based on one type of stabilizer create a bias in observed logical error rates. Strategy v can be generalized to any stabilizer code, and the assignment strategy based only on the support of $X$ or $Z$ stabilizers will increase the error rate of one of $X$ or $Z$ and decrease the other. This result implies that the stabilizer-based assignment may be useful in quantum error correction codes with different $X$ and $Z$ distances.
Both Fig. 7 (A) and (B) showed that the strategy iv randomness + threshold outperformed the other strategies. Maximizing the distance between qubits while also introducing randomness gives the largest boost in performance against logical errors. Note that no assignment strategy does better than the case with $m=1$ where no multiplexing is used.
@@ 5.1 Hypergraph Product Code Structure
$H_{Z}$ $=$ $(I_{n_{1}}\otimes H_{2}|H_{1}^{T}\otimes I_{r_{2}}).$ (4)
These matrices satisfy the condition $H_{X}H_{Z}^{T}=0$ by construction and hence define a valid CSS code $\text{HGP}(H_{1},H_{2})$. When $H_{1}$ and $H_{2}$ are low-density parity checks (LDPC), $H_{X}$ and $H_{Z}$ will also be LDPC. The sizes of $H_{X}$ and $H_{Z}$ are determined by the sizes of the input classical matrices according to the formulas
@@
HGP codes have a geometrically rich Tanner graph structure which can be visualized as the cartesian product of the Tanner graphs for the two input classical codes as shown in Fig. 9. The subgraph corresponding to each row and column in this Tanner graph block structure can be understood as the classical Tanner graph for one of the classical codes used in the construction. As shown in the figure, qubits are represented by circular nodes, and stabilizer checks of both types are represented by square nodes. Additional details regarding this construction are discussed in Appendix D.1.
Surface codes may also be recovered as a special case of hypergraph product code. Using parity check matrices $H_{1}$ and $H_{2}$ for a classical repetition code, $\text{HGP}(H_{1},H_{2})$ is exactly the toric code. Hence, adapting the multiplexing strategies discussed in Sec. 4 to this more general class of codes is a natural next question. However, the linear-time maximum-likelihood generalization of the peeling decoder [55] used in our previous simulations is only defined for the special case of the surface code. This decoder leverages the lattice structure of the surface code to ensure that the erasure subgraph can be completely peeled (Appendix C.3), but this technique does not apply to generic HGP codes. Instead, we introduce another generalization of the peeling algorithm to extend our numerical analysis to HGP codes as well.
@@ 5.2 Pruned-Peeling + VH Decoder
The pruned peeling + VH decoder [59] is a generalization of the peeling algorithm (outlined in Appendix C) specifically designed for HGP codes. It has quadratic complexity and close to maximum-likelihood performance at low erasure rate, making it practically useful for our simulations. This decoder is a modified version of the standard classical peeling decoder based on analysis and correction of two common types of stopping sets, which are patterns of erased qubits that cannot be corrected by simple peeling.
A stabilizer stopping set occurs when the erasure pattern covers the qubit support of an $X$- or $Z$-type stabilizer. The pruned peeling decoder attempts to fix these by removing a qubit from the erasure, thus "breaking" the stabilizer support and possibly allowing the peeling algorithm to become unstuck. Classical stopping sets are patterns of erased qubits supported entirely on a single row or column in the HGP Tanner graph block structure of Fig. 9; any stopping set for a HGP code can be decomposed into a union of components of this form. The VH decoder algorithm attempts to order and efficiently solve each of these classical stopping sets in sequence. The combination of these decoding strategies is referred to as the combined decoder (peeling + pruned peeling + VH). A more detailed explanation is included in Appendix C.4.
The combined decoder is not a maximum likelihood decoder. In addition to logical errors, there still exist patterns of erased qubits where the decoder becomes stuck in a stopping set, leading to a decoder failure. We introduce the term error recovery failure to refer to both decoder failures and logical errors (discussed in Appendix D.2). Our numerical simulations for HGP codes are always with respect to the error recovery failure rate.
@@
Figure 10: Performance of non-ML combined decoder (peeling + pruned peeling + VH) as shown in solid curves and the surface code peeling decoder as shown in dashed curves applied to the $[\![200,2,10]\!]$ toric code using the uniformly random assignment strategy with different numbers of qubits in a single photon, $m$.
@@ 5.3 Assignment Strategies for HGP Codes
As with the surface code, quantum multiplexing can also be utilized with HGP codes. In this section, we analyze the performance of HGP code communication in scenario (C). The scenarios previously proposed in Sec. 4 are also valid for HGP codes, but unlike the special case of the surface code, the distance between any two qubits in a generic HGP code is not easily inferred from a grid. Hence, we do not consider the previously introduced strategies which use distance. We also introduce several new strategies for HGP codes based on stopping sets for the pruned peeling + VH decoder. These strategies are summarized in Table 2, and technical details are included in Appendix D.3.
@@
Strategy i: random The simplest assignment strategy is based on assigning qubits to photons at random.
@@ 6 Discussion and Conclusion
For multiplexed quantum communication, if multiple qubits in a single code word are encoded into the same photon, a correlation of errors in those qubits will be introduced. The simulation results show that it leads to an increase in the logical error rate. We showed that this performance gap can be significantly mitigated by introducing a code-aware (or decoder-aware) strategy to assign qubits to photons, which exploits code structure. In particular, for the surface codes, randomness and also distance maximization are important factors for achieving this. For HGP codes with the VH decoder, minimizing decoder failures was found to be the most important factor.
These techniques can also be exploited to benefit other families of codes and decoders. Furthermore, it is possible to deal with the gap by increasing the code size. We have also shown that it is possible to introduce biased error by using a stabilizer-based assignment strategy. In the special case of the diagonal strategy for the HGP code of Fig. 11, we see that the photon-correlated errors offer an improvement over the no-multiplexing case. In this example, the improvement can be explained by the fact that the diagonal strategy reduces logical errors in addition to decoder failures. Furthermore, this shows the existence of strategies that improve over no-multiplexing despite the fact that fewer resources are used.
@@ Acknowledgements
SN acknowledges Dan Browne, Antonio deMarti iOlius, and Hon Wai Lau for valuable discussions throughout this project. This work was supported by JSPS KAKENHI Grant Number JP21H04880, JP22J20882, the MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant Number JPMXS0118069605, the JST Moonshot R&D Grant Number JPMJMS2061 and JPMJMS226C, and a travel budget from the National Institute of Informatics.
Appendix A Encoding $2^{k}$ dimensional quantum information in a photon
@@
Figure 14: Level $i$ block used in Fig. 13. It split single mode into two modes with new component of time-bin DOF introduced to the state.
Here we show the circuit for encoding $2^{k}$ dimensional quantum information into time-bin in one photon in Fig. 13 where level $i$ block is defined in Fig. 14. In the first half of the circuit, level $i$ block is used to introduce a new component of the time-bin by applying a delay. Then, in the middle of the circuit, each state corresponds to a different mode. In the second half of the circuit, multiplexed photons are output to one mode by applying the optical switches, forming a complete binary tree. This encoding circuit can prepare a photon with $2^{k}$ dimensional time-bin and $2$ dimensional polarization in linear time for $k$.
Appendix B Assignment strategies for surface codes
In this appendix, we describe the details and discussions on each assignment strategy.

### v2

@@ 4.1 Assignment Strategies for Surface Codes
In this section, we describe five strategies for assigning qubits to photons which take advantage of multiplexing, and then evaluate the impact of these strategies on communication performance. These strategies assume that each photon contains a fixed number of qubits $m$ and can be applied in both scenarios (B) and (C). Because (B) can be regarded as a scaled version of scenario (C), we will focus specifically on surface code communication using the latter case, where we send the original code word with $\lfloor 2d^{2}/m\rfloor$ photons.
Strategy i and ii: pair with minimum and maximum distance Strategies i and ii are applicable to the case with $m=2$. Strategy i assigns the nearest neighbor pair of qubits to the same photon, which forms an L-shape in the 2D lattice of the surface code as shown in the example of Fig. 5(A). This minimizes the Manhattan distance in the lattice between the qubits in the same photon. Strategy ii assigns the qubit at coordinates $(i,j)$ on the lattice and the qubit at coordinates $(i+d/2-1\,\textrm{mod}\,d,j+d/2-1\,\textrm{mod}\,d)$ to the same photon. An example is shown in Fig. 5(B). This is the arrangement that maximizes the Manhattan distance, in contrast to strategy i.
Figure 5: Examples showing possible assignments of qubits to photons. Each colored circle denotes a qubit, and the color indicates the photon to which the qubit is assigned. Each non-colored (gray) small circle denotes auxiliary qubits for stabilizer measurement. Strategy i (shown in (A)) minimizes the distance between qubits in the same photon, while strategy ii (shown in (B)) maximizes this distance. Note that this code is defined on the torus represented as a lattice with periodic boundary conditions.
Strategy iii: random Strategy iii is a method in which qubits are selected uniformly at random and assigned to photons.
Strategy iv: random + threshold Strategy iv is designed to increase the separation between qubits assigned to the same photon while exploiting randomness. The strategy works by randomly selecting qubits and accepting them as the set for a photon only if the distance is greater than a certain threshold. If no suitable set of qubits can be found, then the threshold value is reduced.
@@
We describe the details and the motivation of each strategy in Appendix C.
@@
Figure 7: Comparison of multiplexing photon-assignment strategies for surface codes. Logical $Z$ error rate versus photon loss probability. The black curve shows the case without multiplexing. (A) The code parameters are $[\![200,2,10]\!]$ and $m=2$. The gray/brown curve shows the case for the assignment strategy for minimizing (strategy i) / maximizing (strategy ii) the distance between a pair of qubits in the same photon. The orange curve shows the case for uniform randomness (strategy iii), and the blue line shows random + threshold (strategy iv), based on Algorithm 1. (B) The code parameters are $[\![288,2,12]\!]$ and $m=4$. The $Z$ stabilizer-based assignment (light blue curve) outperformed the $X$ stabilizer-based assignment (light orange curve) for logical $Z$ errors. The mixed stabilizer-based assignment strategy performs between $X$ and $Z$. Strategy iii (orange circle) and iv (blue square) outperform the other assignment strategies when the photon loss probability is low.
@@ 4.2 Performance of the assignment strategies
Fig. 7 (A) shows the performance of strategies i through iv, which are applicable to the case of $m=2$. The distance-maximizing strategy (grey) outperforms the distance-minimizing strategy (brown). Logical errors in the surface code correspond to errors covering a longitude or meridian curve on the torus (a vertical or horizontal closed loop in the periodic lattice). When decoding erasure errors, logical errors can only occur when the qubit-support of one of these vertical or horizontal loops is entirely erased. When adjacent qubits in the lattice are erased, as in the case with the distance-minimizing photon assignment strategy, clusters of errors are more likely to cover such loops in the torus. Hence, it is not surprising that the distance-maximizing strategy outperforms the distance-minimizing strategy in our numerical simulations. We also see that the strategies with randomness (iii and iv) outperform the deterministic ones (i and ii). In particular, strategy iv outperformed the other strategies, although there was an increase in logical $Z$-error probability compared to no multiplexing.
Next, we compare the logical $Z$ error rates of strategies iii, iv, and v with $m=4$ in Fig. 7 (B). Assignment strategies based on one type of stabilizer create a bias in observed logical error rates. Strategy v can be generalized to any stabilizer code, and the assignment strategy based only on the support of $X$ or $Z$ stabilizers will increase the error rate of one of $X$ or $Z$ and decrease the other. This result implies that stabilizer-based assignment may be useful in quantum error correction codes with different $X$ and $Z$ distances.
Both Fig. 7 (A) and (B) showed that the strategy iv random + threshold outperformed the other strategies. Maximizing the distance between qubits while also introducing randomness gives the largest boost in performance against the occurrence of logical errors. Note that no assignment strategy does better than the case with $m=1$ where no multiplexing is used.
@@
Refer to caption Refer to caption
(A) (B)
Figure 9: (A) Comparison of $m=2$ photon-assignment strategies on the $[\![200,2,10]\!]$ surface code with the combined error model. The combined error consists of erasure errors and $1$% Pauli $Z$ errors occurring before and after the multiplexed communication. (B) The logical error rate for $m=1$ and $m=2$ (strategy (iv)) with different probability of Pauli $Z$ errors $P_{Z}$. $10^{6}$ samples are taken. Color indicates $P_{Z}$, and solid and dashed lines indicate $m=1$ and $m=2$, respectively.
To assess the impact of multiplexing on the logical error rate in more realistic use cases, we evaluated the performance in a mixed error model where both Pauli errors and photon losses are present. We performed a simulation with the following flow.
(1) Random Pauli errors on sender’s processor $\xrightarrow{\rm transfer\,out\,and\,send}$ (2) Erasure errors on (multiplexed) photonic states $\xrightarrow{\rm transfer\,in}$ (3) Random Pauli errors on receiver’s processor $\rightarrow$ (4) Quantum nondemolition measurement to detect erasure errors $\rightarrow$ (5) Replace erased qubits with mixed state $\rightarrow$ (6) Stabilizer measurement which maps erasure errors to random Pauli errors $\rightarrow$ (7) Decoding procedure. Here, “transfer out” is a procedure to encode qubits into a photon, and “transfer in” is the opposite procedure.
Note that random Pauli errors in (1) have a smaller effect on the logical error rate than in (3) because some erasure errors in (2) will eliminate the effect of (1) on the erased qubits. In this setting, there are Pauli errors only before and after the communication part, and there are no Pauli errors on the multiplexed channel. The behavior of Pauli errors on the multiplexed state is not as obvious as the single-qubit Pauli errors because, for such a case, the Pauli errors may be on qudits, and it can have a larger effect than single-qubit Pauli errors. This qudit Pauli on the multiplexed channel is complex, so we will leave this problem to future work.
Note that we have implemented and used the Union-find decoder [58] for this new simulation, which can handle the combined errors. Fig. 9 (A) shows the logical error rate with the combined error for different strategies. The performance of strategies (i) and (ii) is reversed compared to the case without Pauli error in Fig. 7 (A). The strategy (iv) achieves a logical error rate relatively closer to $m=1$ than (iii) does. Fig. 9 (B) shows the logical error rate versus photon loss probability with different Pauli error rates $P_{Z}$ for $m=1$ and $m=2$ (strategy (iv)). The $m=1$ case outperforms strategy (iv) for low values of $P_{Z}$, but for high values, strategy (iv) is close to no multiplexing and outperforms only slightly for some loss error rates. These indicate that the method of choosing an appropriate assignment strategy in more practical cases may require further consideration.
Note that the threshold [53] for the surface code with Union-find decoder is $9.9$ [58], but here we have random Pauli $Z$ errors occurring before and after the communication process, and so we expect the threshold to be roughly half the original threshold.
@@ 5.1 Hypergraph Product Code Structure
$H_{Z}$ $=$ $(I_{n_{1}}\otimes H_{2}|H_{1}^{T}\otimes I_{r_{2}}),$ (4)
where $I_{n_{1}}$, $I_{n_{2}}$, $I_{r_{1}}$, $I_{r_{2}}$ are identity matrices with sizes $n_{1}$, $n_{2}$, $r_{1}$, and $r_{2}$, respectively. These matrices satisfy the condition $H_{X}H_{Z}^{T}=0$ by construction and hence define a valid CSS code $\text{HGP}(H_{1},H_{2})$. When $H_{1}$ and $H_{2}$ are low-density parity checks (LDPC), $H_{X}$ and $H_{Z}$ will also be LDPC. The sizes of $H_{X}$ and $H_{Z}$ are determined by the sizes of the input classical matrices according to the formulas
@@
HGP codes have a geometrically rich Tanner graph structure which can be visualized as the cartesian product of the Tanner graphs for the two input classical codes as shown in Fig. 10. The subgraph corresponding to each row and column in this Tanner graph block structure can be understood as the classical Tanner graph for one of the classical codes used in the construction. As shown in the figure, qubits are represented by circular nodes, and stabilizer checks of both types are represented by square nodes. Additional details regarding this construction are discussed in Appendix E.1.
Surface codes may also be recovered as a special case of hypergraph product code. Using parity check matrices $H_{1}$ and $H_{2}$ for a classical repetition code, $\text{HGP}(H_{1},H_{2})$ is exactly the toric surface code. Hence, adapting the multiplexing strategies discussed in Sec. 4 to this more general class of codes is a natural next question. However, the linear-time maximum-likelihood generalization of the peeling decoder [55] used in our previous simulations is only defined for the special case of the surface code. This decoder leverages the lattice structure of the surface code to ensure that the erasure subgraph can be completely peeled (Appendix D.3), but this technique does not apply to generic HGP codes. Instead, we introduce another generalization of the peeling algorithm to extend our numerical analysis to HGP codes as well.
@@ 5.2 Pruned-Peeling + VH Decoder
The pruned peeling + VH decoder [60] is a generalization of the peeling algorithm (outlined in Appendix D) specifically designed for HGP codes, where $VH$ stands for Vertical-Horizontal. It has quadratic complexity and close to maximum-likelihood performance at low erasure rate, making it practically useful for our simulations. This decoder is a modified version of the standard classical peeling decoder based on analysis and correction of two common types of stopping sets, which are patterns of erased qubits that cannot be corrected by simple peeling.
A stabilizer stopping set occurs when the erasure pattern covers the qubit support of an $X$- or $Z$-type stabilizer. The pruned peeling decoder attempts to fix these by removing a qubit from the erasure, thus "breaking" the stabilizer support and possibly allowing the peeling algorithm to become unstuck. Classical stopping sets are patterns of erased qubits supported entirely on a single row or column in the HGP Tanner graph block structure of Fig. 10; any stopping set for a HGP code can be decomposed into a union of components of this form. The VH decoder algorithm attempts to order and efficiently solve each of these classical stopping sets in sequence. The name “vertical-horizontal” refers to the fact that, in the HGP Tanner graph picture, individual classical stopping sets are either entirely supported on a column (vertical) or a row (horizontal). The combination of these decoding strategies is referred to as the combined decoder (peeling + pruned peeling + VH). A more detailed explanation is included in Appendix D.4.
The combined decoder is not a maximum likelihood decoder. In addition to logical errors, there still exist patterns of erased qubits where the decoder becomes stuck in a stopping set, leading to a decoder failure. We introduce the term error recovery failure to refer to both decoder failures and logical errors (discussed in Appendix E.2). Our numerical simulations for HGP codes are always with respect to the error recovery failure rate.
@@
Figure 11: Performance of non-ML combined decoder (peeling + pruned peeling + VH) as shown in solid curves and the surface code peeling decoder as shown in dashed curves applied to the $[\![200,2,10]\!]$ surface code using the uniformly random assignment strategy with different numbers of qubits in a single photon, $m$.
@@ 5.3 Assignment Strategies for HGP Codes
As with the surface code, quantum multiplexing can also be utilized with HGP codes. In this section, we analyze the performance of HGP code communication in scenario (C). The scenarios previously proposed in Sec. 4 are also valid for HGP codes, but unlike the special case of the surface code, the distance between any two qubits in a generic HGP code is not easily inferred from a grid. Hence, we do not consider the previously introduced strategies which use distance. We also introduce several new strategies for HGP codes based on stopping sets for the pruned peeling + VH decoder. These strategies are summarized in Table 2, and technical details are included in Appendix E.3.
@@
Strategy i: random The simplest assignment strategy is based on assigning qubits to photons uniformly at random.
@@ 6 Discussion and Conclusion
For multiplexed quantum communication, if multiple qubits in a single code word are encoded into the same photon, a correlation of errors in those qubits will be introduced. The simulation results show that it leads to an increase in the logical error rate. We showed that this performance gap can be significantly mitigated by introducing a code-aware (or decoder-aware) strategy to assign qubits to photons, which exploits code structure. For surface codes in particular, both randomness and distance maximization are important factors for achieving this. For HGP codes with the VH decoder, minimizing decoder failures was found to be the most important factor.
These techniques can also be exploited to benefit other families of codes and decoders. Furthermore, it is possible to deal with the gap between the multiplexing and no-multiplexing cases by increasing the code size. We have also shown that it is possible to introduce biased error by using a stabilizer-based assignment strategy. In the special case of the diagonal strategy for the HGP code of Fig. 12, we see that the photon-correlated errors offer an improvement over the no-multiplexing case. In this example, the improvement can be explained by the fact that the diagonal strategy reduces logical errors in addition to decoder failures. Furthermore, this shows the existence of strategies that improve over no-multiplexing despite the fact that fewer resources are used.
@@ Acknowledgements
SN acknowledges Dan Browne, Antonio deMarti iOlius, and Hon Wai Lau for valuable discussions throughout this project. This work was supported by JSPS KAKENHI Grant Number JP22KJ1436, JP21H04880, the MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant Number JPMXS0118069605, the JST Moonshot R&D Grant Number JPMJMS2061 and JPMJMS226C, and a travel budget from the National Institute of Informatics.
Appendix A Encoding $2^{2}$ dimensional quantum information in a photon
Here we give the transition of states in the circuit shown in Fig. 1 (B) and (C).
A.1 State transition for Fig. 1 (B)
[…]
A.2 State transition for Fig. 1 (C)
[…]
$\xrightarrow{CX_{\textrm{m}_{2},\textrm{p}}}$ $\alpha\ket{H00}+\beta\ket{V01}$
[…]
$\xrightarrow{CX_{\textrm{m}_{1},\textrm{p}}}$ $\alpha\ket{HS00}+\beta\ket{HL01}$
[…]
Then one can measure matter qubits in $X$-basis to obtain the desired state.
Appendix B Encoding $2^{k}$ dimensional quantum information in a photon
@@
Figure 15: Level $i$ block used in Fig. 14. It splits a single mode into two modes with a new component of time-bin DOF introduced to the state.
Here we show the circuit for encoding $2^{k}$ dimensional quantum information into time-bin in one photon in Fig. 14, where the level $i$ block is defined as in Fig. 15. In the first half of the circuit, a level $i$ block is used to introduce a new component of the time-bin by applying a delay. Then, in the middle of the circuit, each state corresponds to a different mode. In the second half of the circuit, multiplexed photons are output to one mode by applying the optical switches, forming a complete binary tree. This encoding circuit can prepare a photon with $2^{k}$ dimensional time-bin and $2$ dimensional polarization in linear time with respect to $k$.
Appendix C Assignment strategies for surface codes
In this appendix, we describe the details and discussions on each assignment strategy for the surface code.